OpenStudy (anonymous):

Use the Sum or Difference Formula Tan -15

myininaya (myininaya):

45-30=15 so 30-45=-15 so expand tan(30-45) using the difference rule for tan

OpenStudy (anonymous):

Sin30cos45-sin39cos45

OpenStudy (mathmale):

First of all, please present the given problem as tan (-15), not as tan -15. Your job is to use the sum or difference for the sine and/or cosine to come up with an expression for tan (-15). But note that there are sum and difference formulas for the tangent function, whcih means that you could by-pass using the since and cosine entirely. Before you proceed, please find and type in the formulas for tan (a+b) and tan (a-b).

OpenStudy (mathmale):

Amb: the expression you've typed in (sin 30 cos 45) - sin 39 cos 45 needs attention. First of all, check it for typos. Secondly, "cos 45" should show up only once, not twice.

OpenStudy (maddie82502):

tan15deg=tan(45-30) tan45-tan30/1+tan45tan30 get the values and substitute you should get 3-3^1/2 / 3+3^1/2 and when you rationalize you should get 2-3^1/2